# Generalizing Scaling Laws for Dense and Sparse Large Language Models

### v2

@@ Abstract
Over the past few years, the size of language models has grown exponentially, as has the computational cost to train these large models. This rapid growth has motivated researchers to develop new techniques aimed at enhancing the efficiency of the training process. Despite these advancements, optimally predicting the model size or allocating optimal resources remains a challenge. Several efforts have addressed the challenge by proposing different scaling laws, but almost all of them are architecture-specific (dense or sparse). In this work we revisit existing scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models. We evaluate and compare our proposed scaling law with existing scaling laws to demonstrate its effectiveness.
Index Terms:
LLMs, Dense models, Sparse models, Scaling laws, Sparsity
@@ I Introduction
In recent years, transformer architectures [1] have revolutionized the deep learning approach. These architectures are now the foundation for the majority of popular large language models (LLMs). As significantly increasing the model size, dataset size and the amount of compute used for training, LLM performance improves dramatically. Because of the large size of LLM models and the large amount of data needed for training, however, the overall training of LLMs becomes time-consuming and expensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training are crucial and have become an important topic in artificial intelligence research. Several approaches have been explored to study the scaling behavior of transformer models to propose scaling laws to determine the optimal allocation of compute, to select optimal model size and data volume for training foundation models efficiently, and to estimate the performance prior to training. Until now, however, nearly all the approaches [2, 3, 4, 5, 6, 7, 8] have been architecture specific (dense or sparse). In this work, we generalize the existing empirical scaling laws and propose a generalized scaling law for both dense and sparse models that enables accurate performance prediction and optimal resource allocation across different architectures.
The field of natural language processing has undergone significant advancement with the development of transformer architectures [1] such as BERT [9] and GPT-3 [10]. Since ChatGPT was publicly released and made available in the fall of 2022, LLMs have gained more attention across diverse domains as a key innovation technology. Several models [11, 12, 13, 14, 15] have been developed over the past few years with traditional dense architecture. A variation of the transformer architecture where a fraction of available parameters are activated—namely, sparse transformers[16]—has become popular lately, and several models [17, 18, 19, 20, 21, 22, 23, 24] have been developed using the concept of sparsification. One common factor among all these models is that as they are evolving: their size have been increasing significantly from GPT-2 with $1.5$ billion parameters [25] to PaLM with 540 billion parameters [14], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [26], which has 94.49% sparsity. With the increase of size, these models require large training datasets, eventually increasing the compute budget [2, 3] to several petaFLOP/s-days of computation to achieve optimal performance, thus making the models expensive to train.
@@
In traditional transformer architectures, parameters of every layer are active for every token. Each token flows through the same feed-forward network; attention layers analyze the entire input sequence simultaneously, and every attention head considers every other position. These kinds of models are dense. The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset. But it comes with high computational cost and scales up with both model size and sequence length.
One way to minimize the cost is to introduce sparsity in a model, where not all the parameters available in the model are used to process a token. Sparsity can be introduced in several ways. We focus on two methods: pruning [27] and mixture of experts [28]. Pruning-based sparsity permanently removes parameters from the model. On the other hand, MoE introduces new experts and uses only a fraction of them to process a token. (A more detailed explanation of pruning and MoE is given in Subsection II-B.)
Foundation models have gained attention across diverse domains because of their high accuracy, reasoning abilities, and generative power. Achieving these capabilities, however, requires vast amounts of training data and substantial compute resources. As a result, optimal allocation of compute, model size, and data volume has become critical to minimizing training costs.
@@
Scaling laws—empirical relationships often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several such empirical scaling laws based on different LLM architectures have been proposed in recent years [2, 3, 4, 5, 6, 7, 8, 29]. Others are based on training precision [30], distillation [31], and quantization [32]. Several studies also have been done on vision [4], audio [33, 34], and mixed-modal language models [35]. This interest in scaling laws for different domains makes them pivotal in deep neural network research.
In this paper we focus on the scaling laws for dense and sparse LLMs. We make the following contributions:
@@
  • We propose a generalized scaling law for both dense and sparse LLMs and discusses the relationship between the proposed scaling law and the existing laws (Section III).
@@ II Scaling Laws for LLMs
Transformer models use a fully connected attention mechanism in which the token is addressed by the other tokens of the sequence. This process, which activates all the parameters in the model, is computationally expensive. The cost of training a large transformer model depends on several factors. Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers, which have shown different scaling behaviors. In the following, we discuss the scaling laws for dense and sparse LLMs.
II-A Dense Scaling Laws
Traditional transformer models are dense in nature. Kaplan et al. [2] studied the behavior of LLMs with a series of experiments and different hyperparameters: model size, model shape, total compute, data volume, and batch size. The authors observed that performance or model loss during pretraining follows a power law relationship with the total number of parameters, the total number of training tokens, and the total compute budget. They established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [2] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. They suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting. They represented training loss as a function of number of parameters and data size and proposed the following empirical scaling law:
@@ II-A Dense Scaling Laws
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [3] extended the work in [2] and concluded that model size and total number of tokens should be increased in proportion to avoid overfitting. In other words, if the total number of parameters or model size is doubled, data volume or total number of tokens should be doubled as well. They represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on their observations, the authors proposed the following scaling law:
@@
This law establishes a power-law relationship between loss $L$, number of parameters $N$, and the data size $D$. Here $e$ captures the entropy of the natural text, which is the minimum loss as both $N$ and $D$ approach infinity. The second term (parameter-dependent term) captures the inverse relationship between model size $N$ and loss. The third term (data-dependent term) captures the impact of the data size $D$ on loss. The constants $a$ and $b$ and the exponents $\alpha$ and $\beta$ are determined empirically through some experiments and fitting the data. Its best-fitted values for the coefficients and exponents are listed in Table II. Equation 2 currently is a well-known empirical scaling law for dense LLMs.
@@
II-B Sparse Scaling Laws
Sparse models use only a subset of the parameters available in the models. As a result, Equation 2 needs to be adjusted and extended. Several works had proposed scaling laws for sparse transformer models [4, 5, 8]. These scaling laws are based on different sparsification methods. In this work we focus on two types of sparsifications: pruning and mixture-of-experts (MoE).
II-B1 Pruned Transformer Models
One way to introduce sparsity in transformer models is pruning, where a number of parameters are removed from the architecture and are not used during the forward pass or backpropagation, hence making no contribution to the training process. Pruning can be categorized into three main groups [27]. In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy. In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned, with the importance based on certain criteria. In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network.
Frantar et al. [4] studied the scaling behavior of models pruned by unstructured pruning. Sparsity $S$ is defined as $S=\frac{\text{Total Parameters - Nonzero Parameters}}{\text{Total Parameters}}$. The experimental results with a total of $48$ models trained showed that sparsity introduced by pruning affects only the parameter-dependent term of Equation 2 and has no impact on the data-dependent term. The loss was fitted as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in the following equation.
@@ II-B1 Pruned Transformer Models
Following the similar approach of [3] to find the optimal values for the coefficients of Equation 3, the best-fitted values for these coefficients are determined empirically through some experiments and fitting the data shown in Table III.
@@
To draw the similarities between Equations 2 and 3, we used the similar symbols in Equation 2 to reformat the coefficients in Equation 3. We used $e$ instead of $c$ to capture the lower bound on the loss that represents the inherent stochasticity of the modeling problem. We replaced $b_{N}$ and $b_{D}$ with $\alpha$ and $\beta$, respectively, and ${a_{D}}^{b_{D}}$ with $b$. The result is the following equation:
@@
Equation 4 helps visualize the transition from the dense scaling law Equation 2 to the sparse scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity. Based on the changes in the coefficient representation, the best-fitted values need to be updated correspondingly. The updated values for the coefficients in Equation 4 are calculated and listed in Table IV.
@@
While Equation 3 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [4]. We have generated an estimated IsoFLOP dataset for the compute budget $1e20$ from the experiments conducted in [5], which included sparsities of more than $87.5\%$. When we used the new dataset to plot the scaling behavior of Equation 3, we noticed a spike for smaller nonzero parameters, as shown in Figure 1, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with nonzero parameters as x-axis and training loss as y-axis. Increasing the nonzero parameters results in decreasing the data size. However, based on the work in [4, 5], sparsity should provide a smooth performance gain as we increase sparsity. This indicates that Equation 3 does not capture the scaling behavior of highly sparse models very well.
Figure 1: IsoFLOP plot of Equation 3 for different sparsities
On the other hand, the largest model used in [4] had only $85M$ nonzero parameters. Compared with the current work, these models are relatively much smaller. This work used the scaling law Equation 2 [3] as the base for the approach. For $0\%$ sparsity, or for dense models, the loss should be equal to Equation 2. But for the same data set with no sparsity, Equation 3 provides different coefficient values from those of Equation 2, as shown in Figure 2, there is a difference of up to 2.5 in loss. Thus, we can conclude that Equation 2 is not a special case of Equation 3 if the sparsity is 0.
Figure 2: Comparison between Equations 2 and 3 for dense models
@@ II-B2 MoE Models
Another way to introduce sparsity is to use the mixture-of-experts method [28] where the feed-forward network for a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token, and this number is a fixed value. The tokens are routed to different experts during the pretraining process. The size of the experts is usually equal to the size of the feed-forward network.
Abnar et al. [5] investigated the sparse MoEs through a large-scale empirical study. The sparsity $S$ was defined as $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts. For five different compute budgets and seven different sparsities, different models were trained. Based on the experimental results, the authors proposed the following scaling law:
@@
Equation 5 introduced two different terms in the scaling law along with the previous three terms from [3] to capture the impact of sparsity. The newly introduced terms are interdependent. The fifth term in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S. Because the total number of parameters already includes the active parameters, however, having two different terms for total parameters and active parameters in Equation 5 causes some redundancy in the loss function.
This work also used Equation 2 as the base. But for $0\%$ sparsity we apply the models from [5] to Equations 5 and 2 to generate Figure 3 and notice a difference in loss values. This indicates that Equation 2 is not a special case of Equation 5 if the sparsity is 0.
Figure 3: Comparison between Equations 2 and 5 for dense models

### v3

@@ Abstract
Despite recent advancements of large language models (LLMs), optimally predicting the model size for LLM pretraining or allocating optimal resources still remains a challenge. Several efforts have addressed the challenge by proposing different empirical scaling laws, but almost all of them are architecture-specific (dense or sparse). In this work we revisit existing empirical scaling laws and propose a generalized scaling law to provide a unified framework that is applicable to both dense and sparse large language models. We evaluate and compare our proposed scaling law with existing scaling laws and demonstrate that our proposed scaling law captures the scaling behavior of existing scaling laws. Further, we show an IsoFLOP comparison between our proposed scaling law and the state-of-the-art scaling law to illustrate the effectiveness of our proposed scaling law for Mixture-of-Expert (MoE)-based very large LLMs like DeepSeek-V3. Our proposed scaling law can be used to estimate the best model hyperparameters (Model size, Tokens and Compute) for a given sparsity or to identify the optimal sparsity for the given model hyperparameters.
@@ I Introduction
In recent years, transformer architectures [37] have revolutionized the deep learning approach especially the field of natural language processing with the development of transformer-based architectures [37] such as BERT [12] and GPT-3 [3] etc. Transformer architecture is now the foundation for the majority of popular large language models (LLMs). Since the release of ChatGPT LLMs have gained more attention across diverse domains as a key innovative technology. With the growing popularity across diverse domain there is an increasing need to scale LLM knowledge and accuracy while also reducing the inference latency and training cost. However, achieving these goals is fundamentally challenging due to the computational structure of standard transformer models, as computation grows rapidly with the increase of model size and context length.
@@
In traditional transformer architecture, parameters in every layer are activated to process every token. Each token passes through the same feed-forward network, and in a multi-head attention setting, every attention considers every other position. These traditional forms of transformers are called dense transformers. The complex and exhaustive architecture allows the dense models to understand the rich contextual meaning of the input dataset, but it comes with a high computational cost that scales with both model size and sequence length.
One way to reduce the cost is to introduce sparsity where all the available parameters are not used to process a token in a sparse model. There are several ways to introduce sparsity, but in this work we focus on two methods: pruning [6] and mixture of experts (MoE)  [33]. Pruning-based sparsity permanently removes parameters from the model. Pruning can be categorized into three main groups [6] namely: structured pruning, unstructured pruning, and semi-structured pruning. In structured pruning, entire higher-level units or even whole layers are removed; this process delivers universal speedup but sometimes at the cost of accuracy. In unstructured pruning, individual weights are removed from anywhere in the network; usually the least important weights are pruned. In semi-structured pruning, one tries to strike a balance between unstructured and structured pruning approaches by enforcing fixed pruning patterns in small regions of the network. On the other hand, in MoE-based models the feed-forward network for a transformer model is replicated multiple times, and each of these copies is called an “expert”. In the MoE-based models only a fraction of experts are activated per token, instead of all the experts and that’s how sparsity is introduced. The number of activated experts is a fixed value and determines the sparsity level. The tokens are routed to different experts during the pretraining process. The size of the experts is usually equal to the size of the feed-forward network but can differ too.
Several models [31, 34, 36, 8, 28] have been developed over the past few years with traditional dense architecture. But lately, sparse transformers[7]—have become extremely popular, and several models [14, 15, 43, 27, 25, 10, 22, 18] have been developed using the concept of sparsification. One common factor among all these models is that as they are evolving, their size has been increasing significantly from GPT-2 with $1.5$ billion parameters [30] to PaLM with 540 billion parameters [8], to DeepSeek-V3, a strong mixture-of-experts (MoE) language model with 671 billion parameters, 37 billion of which are activated for each token [11], which has 94.49% sparsity. The reason is, it has been proven that as model size, dataset size, and the amount of compute used for training increase, LLM performance improves smoothly [23]. However with the increase of size, these models require large training datasets, eventually increasing the compute budget [23, 20] to several petaFLOP/s-days of computation to achieve optimal performance. This makes LLM training time-consuming and resource intensive. As a result, determining the optimal allocation of resources, model size, and training data and estimating the training performance before conducting the actual training is crucial.
@@
Scaling laws are the empirical relationships, often expressed as power laws—describe a model’s performance based on allocated resources. These laws help determine the optimal model size and data volume for a given compute budget to achieve optimal pretraining performance. Several empirical scaling laws based on different LLM architectures have been proposed in recent years [23, 20, 16, 1, 19, 9, 26, 17]. Few scaling law works focused on training precision [24], distillation [4], and quantization [5]. Several studies have also been done on vision [16], audio [13, 29], and mixed-modal language models [2]. This interest in scaling laws for different domains makes them pivotal in deep neural network research. However, until now, nearly all existing approaches have been architecture-specific (dense or sparse) and also use different scaling law coefficients (details in Section II) which makes it difficult to perform cross-architecture comparisons and budget planning. In this work we show that transformers, regardless of dense and sparse architectures, exhibit a common scaling behavior which depends on the number of active/nonzero parameters. To demonstrate that, we generalize the existing empirical scaling laws to propose a scaling law that enables accurate performance predictions and optimal resource allocation across architectures. In this paper we focus on the scaling laws for both dense and sparse LLMs. We make the following contributions:
@@
  • We propose a generalized scaling law for both dense and sparse LLMs and discuss the relationship between the proposed scaling law and the existing laws (Section III).
@@ II Scaling Laws for LLMs
Based on the number of parameters available in the architecture and the number of parameters that are being used during pretraining, transformer models can be categorized into dense and sparse transformers. Previous works [31, 34, 36, 8, 28, 14, 15, 43, 27, 25, 10, 22, 18] have shown that these models exhibit different behavior during pretraining as they scale, and hence their scaling behaviors are studied separately.
II-A Scaling Laws for Dense Models
Traditional transformer models are dense in nature. Kaplan et al. [23] studied the behavior of dense LLMs with a series of experiments by monitoring different hyperparameters: model size, model shape, total compute, data volume, and batch size during pretraining. The authors established that model loss depends strongly on scale, which consists of the number of model parameters $N$, the dataset size (number of tokens) $D$, and the amount of compute $C$ used for training. It also depends mildly on other architectural hyperparameters such as model shape: depth and width. For dense models, the compute $C$ (training FLOPs) required to train a transformer model with $N$ parameters and $D$ tokens is $C=6ND$ [23] because approximately 6 floating point operations are needed for every parameter in the model and every training sample. The authors also suggested that if compute is increased by $10$ times, then model size should be increased by $5.5$ times and dataset size by $1.8$ times to avoid overfitting. They represented training loss as a function of number of parameters and data size and proposed the empirical scaling law:
@@ II-A Scaling Laws for Dense Models
Using three different approaches—fixed model size with varying numbers of training tokens, IsoFLOP profiling, and parametric function fitting—Hoffmann et al.  [20] extended and corrected Equation 1 in [23] and concluded that model size and number of tokens should be increased in proportion to avoid overfitting. The authors represented training loss as a combination of irreducible loss or entropy $e$, a parameter-dependent term or the loss incurred by a model of fixed parameter size, and a data size term or the loss induced by a fixed number of training steps. Based on the observations from their three experiments, the authors proposed Equation 2 as their scaling law.
@@
Figure 1: Comparison between Hoffman scaling law and Frantar scaling law at $0\%$ sparsity.
This law establishes a power-law relationship between loss $L$, NN, and $D$. Here $e$ captures the entropy of the natural text, which is the minimum loss as both $N$ and $D$ approach infinity. The second term (parameter-dependent term) captures the inverse relationship between $N$ and loss. The third term (data-dependent term) captures the impact of $D$ on loss. The constants $a$ and $b$ and the exponents $\alpha$ and $\beta$ are determined empirically through some experiments and fitting the data. Its best-fitted values for the coefficients and exponents are listed in Table II. Equation 2 currently is the state-of-the-art empirical scaling law for dense LLMs. For the rest of the paper, we will refer Equation 2 as Hoffman scaling law.
@@
II-B Scaling Laws for Sparse Models
Sparse LLMs use only a subset of the parameters available in the models. Several works had proposed scaling laws for sparse transformer models [16, 1, 26]. These scaling laws are based on different sparsification methods. In this work we focus on two types of sparsifications: pruning and mixture-of-experts (MoE).
Figure 2: Comparison between Hoffman scaling law and Abnar scaling law at $0\%$ sparsity.
II-B1 Pruned Models
One way to introduce sparsity in transformer models is pruning, where a subset of parameters are removed or masked and are not used during the forward pass or backpropagation, hence making no contribution to the training process. Frantar et al. [16] studied the scaling behavior of models pruned by unstructured pruning. In their work, the authors defined sparsity, $S=\frac{\text{Total Parameters}-\text{Nonzero Parameters}}{\text{Total Parameters}}$. Based on the pretraining results of $48$ models with different sparsity levels, numbers of nonzero parameters and training tokens, the authors claimed that sparsity introduced by pruning affects only the parameter-dependent term of Hoffman scaling law and has minor to no impact on the data-dependent term. The authors also claim that increasing sparsity leads to a better pretraining loss. Their experimental results also showed that as the number of nonzero parameters and training tokens—in other words, total compute increased the pretraining loss keeps getting better, which holds the claim of Hoffman scaling law [20]. Using the Hoffman scaling law as the foundation, the authors fitted loss as a function of nonzero parameters $N$, data $D$, and sparsity $S$ as shown in Equation 3.
@@ II-B1 Pruned Models
Following the similar approach of [20] the authors presented the best-fitted values for the scaling law coefficients of Equation 3. Table III lists the scaling law parameter values mentioned in [16]. For the rest of the paper, we will refer to Equation 3 as the Frantar scaling law.
@@
To draw the similarities between Hoffman scaling law in Equation 2 and Frantar scaling law in Equation 3, we used the similar symbols used in Hoffman scaling law to reformat the scaling law coefficients in Frantar scaling law. We used $e$ instead of $c$ to capture the lower bound on the loss that represents the inherent stochasticity of the modeling problem. We replaced $b_{N}$ and $b_{D}$ with $\alpha$ and $\beta$, respectively, and ${a_{D}}^{b_{D}}$ with $b$. $c_{S}$ is the constant sparsity factor. The reformatted version of Frantar scaling law is Equation 4.
@@
Equation 4 helps visualize the transition from the dense Hoffman scaling law Equation 2 to the sparse Frantar scaling law Equation 3 in order to investigate which factor is directly impacted by sparsity. Based on the changes in representation, we recalculate the best-fitted values for the scaling law coefficients in the reformatted Frantar scaling law, and the values are listed in Table IV.
@@
While Equation 4 functionally captures the effect of sparsity on loss, the maximum sparsity used in the experiments was $87.5\%$ [16] while other works [1] show that better performance gain can be achieved by introducing even more sparsity (even at 98% sparsity). The largest model used in Frantar scaling law experiments had only $85M$ nonzero parameters which are relatively much smaller. Although this work used the Hoffman scaling law as baseline, at $0\%$ sparsity or for dense models with the same dataset, Equation 4 provides different loss predictions than those in Hoffman scaling law with an average MSE (Mean Squared Error) of $3.04$, as shown in Figure 1. Thus, we can conclude that Hoffman scaling law is not a special case of Frantar scaling law where sparsity is $0\%$
Figure 3: Loss prediction of Hoffman scaling law and the proposed scaling law.
@@ II-B2 MoE Models
Another way to introduce sparsity is to use the mixture-of-experts (MoE) method [33] where the feed-forward network of a transformer model is replicated multiple times and each of these copies is called an “expert.” Only a fraction of experts are activated per token. The tokens are routed to different experts during the pretraining process.
(a) Training tokens: $16B$
(b) Training tokens: $32B$
(c) Training tokens: $65B$
Figure 4: As sparsity and the number of nonzero/active parameters increase, the pretraining loss decreases. In each figure, the number of non-zero parameters in the models used ranged from 1.3M85M1.3M-85M with varying sparsity levels, and the number of training tokens is mentioned in the subcaption.
Abnar et al. [1] investigated the sparse MoEs through a large-scale empirical study. The authors defined sparsity, $S=\frac{E-K}{E}$, where $E$ is the total number of experts and $K$ is the number of active experts. For five ($3e19,6e19,1e20,3e20,\text{ and }1e21$) different compute budgets and seven ($0\%,25\%,50\%,75\%,90\%,95\%,\text{ and }98\%$) different sparsities, different models were trained. The conclusion they made was that as we increase sparsity along with model size and number of training tokens, the pretraining loss decreases. For IsoFLOP experiments, they noticed that each for a specific sparsity level, each model size has a sweet spot. Until that spot, the pretraining loss keeps decreasing, and beyond that spot, pretraining loss keeps increasing. Based on the experimental results, the authors proposed Equation 5 as their scaling law. For the rest of the paper, we will refer to Equation 5 as Abnar scaling law.
@@
Abnar scaling law introduced two different terms in the scaling law along with the previous three terms from Equation 2 [20] to capture the impact of sparsity. The newly introduced terms are interdependent. The fifth term $\frac{d}{(1-S)^{\delta}N^{\gamma}}$ in Equation 5 can be interpreted as an active parameter term because the denominator term is a multiplication of $N$ and 1S1-S. As the total number of parameters already includes the active parameters, having two different terms for total parameters and active parameters in Equation 5 can be deemed as redundant. While this may help the Abnar scaling law to fit the experimental results better, we later show that this can be represented in a generalized way without introducing redundancy.
Abnar scaling law also used Hoffman scaling law as the base. We compared both of these laws at $0\%$ sparsity using the same model size and data size used in Abnar scaling law experiments. The results were off from each other with and average MSE of $1.06$ in Figure 2.